OpenStudy (anonymous):

Confirm that f and g are inverses by showing that f(g(x)) = x and g(f(x)) = x. f(x) =x+4/6 and g(x) = 6x - 4

OpenStudy (zzr0ck3r):

f(g(x)) = f(6x-4) = (6x-4+4)/6 = x = 6((x+4)/6 - 4 = g((x+4)/6) = g(f(x))